Unit Costs, Inventory Valuation, Variable and Absorption Costing

Snyder Company produced 80,900 units during its first year of operations and sold 77,100 at $21.60 per unit. The company chose practical activity—at 80,900 units—to compute its predetermined overhead rate. Manufacturing costs are as follows:

Direct materials $457,085
Direct labor 87,372
Expected and actual variable overhead 300,139
Expected and actual fixed overhead 432,006

Required:

If required, round unit cost answers to the nearest cent.

1. Calculate the unit cost and the cost of finished goods inventory under absorption costing.

Unit Cost $
Cost of finished goods inventory $

2. Calculate the unit cost and the cost of finished goods inventory under variable costing.

Unit Cost $
Cost of finished goods inventory $

3. What is the dollar amount that would be used to report the cost of finished goods inventory to external parties?
$

Why?
Answer #1

Solution 1:

Cost of goods manufactured under absorption costing = Direct material + Direct labor + Variable manufacturing overhead + Fixed manufacturing overhead

= $457,085 + $87,372 + $300,139 + $432,006 = $1,276,602

unit cost = Cost of goods manufactured / Nos of units manufactured = $1,276,602 / 80900 = $15.78 per unit

Cost of finished goods inventory = Ending inventory of finished goods * Cost per unit

= (80900 - 77100) * $15.78 = $59,964

Solution 2:

Cost of goods manufactured under variable costing = Direct material + Direct labor + Variable manufacturing overhead

= $457,085 + $87,372 + $300,139 = $844,596

unit cost = Cost of goods manufactured / Nos of units manufactured = $844,596 / 80900 = $10.44 per unit

Cost of finished goods inventory = Ending inventory of finished goods * Cost per unit

= (80900 - 77100) * $10.44 = $39,672

Solution 3:

Dollar amount that would be used to report the cost of finished goods inventory to external parties = $59,964 because, financial statements are prepared using absorption costing.